# Étale motives of geometric origin

### v1

@@ Abstract
Consider a scheme $X$ and a commutative ring of coefficients $\Lambda$. We denote by $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$ the stable $\infty$-category of étale motives defined as the Tate-stabilisation of the category of $\mathbb{A}^{1}$-local étale hypersheaves of $\Lambda$-modules over the category of smooth $S$-schemes. This version of étale motives without transferts was studied by Ayoub in [Ayo14]; it is equivalent over noetherian schemes to the category of h-motives studied by Cisinski and Déglise in [CD16]. It is endowed with the six functors formalism by [AGV22, Theorem 4.6.1]. Denote by $\Lambda_{X}$ the unit object of $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$.
@@
Our main goal is to prove that motives of geometric origin form an étale hypersheaf and to give a purely categorical characterisation of geometric étale motives:
@@
(4.1 and 2.4) Let $\Lambda$ be a commutative ring.
@@
This question dates back to [CD16], where Cisinski and Déglise considered étale-locally geometric [^1] Beware that in loc. cit., they call constructible what we chose to call of geometric origin in this paper. We chose this terminology because they are the motives that can be built from objects coming from geometry. In addition, objects that are called locally constructible by Cisinski and Déglise are more similar to classical constructible sheaves. In any case, the purpose of this paper is to show that all those notions coincide. motives and showed that over noetherian schemes, the latter coincides with constructible motives. To prove the above theorem, we need to generalise to non-noetherian schemes the continuity properties proved in loc. cit.
@@
(Continuity; 3.2) Let $\Lambda$ be a commutative ring. Consider a qcqs finite-dimensional scheme $X$ that is the limit of a projective system of qcqs finite-dimensional schemes $(X_{i})$ with affine transition maps. Then, the natural map
@@
Using continuity, the most critical case of the main theorem is the case where $X=\mathrm{Spec}\ (k)$ with $k$ a field and where $\Lambda$ is regular and flat over $\mathbb{Z}$. In that particular case, the result can be deduced from the case of rational coefficients which was proved in [CD16] and results of [Rui22] on Artin motives. The latter are consequences of results of [BG23] which imply that étale sheaves of the form $f_{\sharp}\Lambda_{L}$ for $L/k$ a finite separable extension, generate the category of étale-locally perfect constant sheaves as a thick subcategory of itself.
@@
(5.2 and 5.3) Let $\Lambda$ be a commutative ring. The functor $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$ satisfies Milnor excision and h-descent over qcqs finite-dimensional schemes.
@@
Let $X$ be a scheme and let $\Lambda$ be a commutative ring. We denote by $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$ the subcategory of $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$ made of constructible motives, that is those motives such that for any open affine $U\subseteq X$, there is a finite stratification $U_{i}\subseteq U$ made of constructible locally closed subschemes such that each $M|_{U_{i}}$ is dualisable.
@@
In the course of the proof, we will need the following lemma.
Lemma 3.3.
Let $\Lambda$ be a commutative ring. Assume that 3.2 holds with coefficients $\Lambda$. Then, geometric motives with coefficients $\Lambda$ are constructible.
@@
Let $X$ a qcqs scheme. Write $X$ as a limit of noetherian finite-dimensional schemes $X_{i}$ with affine transition maps. Any geometric motive over $X_{i}$ is constructible using [CD16, Theorem 6.3.26]. If $M$ is a geometric object, it is the pullback of some $M_{i}$ over some $X_{i}$ because we assumed 3.2 to be true. As the motive $M_{i}$ is constructible, so is $M$. ∎
Proof.
(of 3.2) Once full faithfulness is proven, the essential surjectivity is easy: we first reduce by Zariski descent to objects of the form $f_{\sharp}(\Lambda_{Y})(n)$ with $Y$ affine and $f\colon Y\to X$ smooth. The usual continuity properties of [Gro66, Théorème 8.10.5, Proposition 17.7.8] then give the result. Therefore we tackle the full faithfulness.
Let $(M_{i})$ and $(N_{i})$ be cartesian sections over the diagram of schemes $(X_{i})$ of the functor $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$, and let $M$ and $N$ be their respective pullbacks along the maps $X\to X_{i}$, we want to prove that the canonical map
$\mathrm{colim}\hskip 2.84526pt\mathrm{Map}_{\mathrm{DM}^{\mathrm{\acute{e}t}}}(M_{i},N_{i})\to\mathrm{Map}_{\mathrm{DM}^{\mathrm{\acute{e}t}}}(M,N)$
is an equivalence.
First assume that $\Lambda$ is flat over the integers (so that $\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z}$ is a non-derived ring for all prime $p$). The above equivalence be tested after tensoring with $\mathbb{Q}$ and $\mathbb{Z}/p\mathbb{Z}$ for any prime $p$. Tensoring with $\mathbb{Z}/p\mathbb{Z}$ is a finite limit and therefore gets inside mapping spectra. Using 3.1 this is also the case for tensoring with $\mathbb{Q}$, thus we can assume that either $\Lambda$ is a $\mathbb{Q}$-algebra, or that $\Lambda$ is a $\mathbb{Z}/p\mathbb{Z}$-algebra. The case of $\mathbb{Q}$-algebras is already known by [EHIK21b, Lemma 2.2 (i), Lemma 5.1 (ii)]. Using the classical argument with the Artin-Schreier exact sequence given in [CD16, Proposition A.3.4], we can assume $p$ to be invertible in $X$ and then use the rigidity theorem [BH21, Theorem 3.1], so that it suffices to prove the statement for constructible étale sheaves with torsion coefficients which was proved in [HS23, Theorem 2.2].
Assume now that we know the result for a ring $R$, we show that the result still holds for $\Lambda=R/(f)$ for any element $f$ of $R$. Let 00 be an element of our index set, so that $M_{i}$ (resp. $N_{i}$) is the pullback to $X_{i}$ of $M_{0}$ (resp. $N_{0}$) for $i\geqslant 0$. To prove the result, by dévissage, we may assume that $M_{0}$ and $N_{0}$ are generators. In particular, they are isomorphic to $M_{0}^{\prime}/(f)$ and $N_{0}^{\prime}/(f)$ for some generators $M_{0}^{\prime}$ and $N_{0}^{\prime}$ of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X_{0},R)$. But for those the mapping spectra is the cofiber of the multiplication by $f$, so that we obtain the result.
Hence, the result is true when $\Lambda=A/J$, where $A$ is flat over $\mathbb{Z}$ and $J$ is a finitely generated ideal of $A$. In particular, the result is true for any ring of finite type over $\mathbb{Z}$.
In general, we may write $\Lambda=\mathrm{colim}_{i}\hskip 2.84526pt\Lambda_{i}$ as a filtered colimit of rings of finite type over the integers and we have the following 3.4.
Lemma 3.4.
Let $\Lambda=\mathrm{colim}_{j\in J}\hskip 2.84526pt\Lambda_{j}$ be a colimit of commutative rings and let $X$ be a qcqs scheme. Then the natural functor
$\mathrm{colim}_{j}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda_{j})\to\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda)$
is an equivalence.
Proof.
This proof was communicated to the authors by Denis-Charles Cisinski. First, we prove that the functor is fully faithful. Let $(M_{j})$ and $(N_{j})$ be objects of the colimit, and let 00 be an element of the index set, so that $M_{j}$ and $N_{j}$ are defined over $\Lambda_{0}$ for $j\geqslant 0$. By dévissage we may assume that $M_{0}$ and $N_{0}$ are generators, so that they are in fact of the form $M\otimes_{\mathbb{Z}}\Lambda_{0}$ and $N\otimes_{\mathbb{Z}}\Lambda_{0}$ with $M$ and $N$ in $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\mathbb{Z})$. Thus (using the $\otimes_{\mathbb{Z}}\Lambda$ adjunctions) we want to prove that the map
$\mathrm{colim}_{j}\hskip 2.84526pt\mathrm{Map}_{\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\mathbb{Z})}(M,N\otimes_{\mathbb{Z}}\Lambda_{j})\to\mathrm{Map}_{\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\mathbb{Z})}(M,N\otimes_{\mathbb{Z}}\Lambda)$
is an equivalence. This can be checked after tensoring with $\mathbb{Q}$ and with $\mathbb{Z}/p\mathbb{Z}$ for all primes $p$. The map, when tensored by $\mathbb{Q}$ is an equivalence because by 3.1 it gives the analogous map with the integers replaced by $\mathbb{Q}$, and the latter is an equivalence because $M\otimes\mathbb{Q}$ is then a compact object of $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\mathbb{Q})$ by [EHIK21b, Lemma 5.1(i)]. After tensorisation by $\mathbb{Z}/p\mathbb{Z}$ (which is a finite colimit, hence commutes with the colimit and the $\mathrm{Map}$), we obtain by rigidity the analogous map in $\mathrm{D}(X_{\mathrm{\acute{e}t}},\mathbb{Z}/p\mathbb{Z})$:
$\mathrm{colim}_{j}\hskip 2.84526pt\mathrm{Map}_{\mathrm{D}(X_{\mathrm{\acute{e}t}},\mathbb{Z}/p\mathbb{Z})}(M\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z},N\otimes_{\mathbb{Z}}\Lambda_{j}\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z})\to\mathrm{Map}_{\mathrm{D}(X_{\mathrm{\acute{e}t}},\mathbb{Z}/p\mathbb{Z})}(M\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z},N\otimes_{\mathbb{Z}}\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z}).$
Now, each complex $\Lambda_{j}\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z}$ is concentrated in degrees $[-1,0]$ hence we are looking at a colimit which is uniformly bounded, and $M\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z}$ is constructible by 3.3 applied to $\Lambda=\mathbb{Z}$. As filtered colimits of uniformly bounded objects commute with mapping spectra from constructible complexes by [HRS23, Lemma 5.3], the full faithfulness holds. The essential surjectivity is easy, as it is clear that generators of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda)$ are in the image. ∎
Finally, we have that
$\mathrm{colim}_{i}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X_{i},\Lambda)\simeq\mathrm{colim}_{i}\mathrm{colim}_{j}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X_{i},\Lambda_{j})\simeq\mathrm{colim}_{j}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda_{j})\simeq\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda),$
finishing the proof of the theorem. ∎
@@
For the first point, assume first that $f=p$ is proper. Let $M$ be an object of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(Y,\Lambda)$, and write $X=\lim X_{i}$ as a limit of schemes that are of finite type over the integers. By [Gro66, Théorème 8.10.5.(xii)] one can find a $p_{i_{0}}\colon Y_{i_{0}}\to X_{i_{0}}$ proper such that the pullback to $X$ is $p$. In particular
@@
such that $F(j)$ is of the form $p^{j}_{*}\Lambda(n_{j})$ for $p^{j}$ projective and $n_{j}\in\mathbb{Z}$, and such that $M_{i}$ is a direct factor of $\mathrm{colim}_{J}F$. Thus, the object $M$ is a direct factor of $(\mathrm{colim}_{J}F)_{\mid X}$, which by proper base change is the colimit of the objects $q^{j}_{*}\Lambda(n_{j})$, with $q^{j}$ the base change of $p^{j}$ to $X$. This finishes the proof, as the third point readily follows from the second one. ∎
Combining 3.2 and 3.3, we proved:
@@
Geometric motives are constructible.
We now prove the converse when $X$ is $\Lambda$-finite by linking both notions to compactness:
@@
We proved in 3.6 that motives of geometric origin are constructible. We now prove the converse. Let $M$ be constructible over $X$. We can assume that $X$ is affine by Zariski hyperdescent of geometric motives. By dévissage and localisation, we can further assume that $M$ is dualisable. In that case $\underline{\mathrm{Hom}}(M,-)$ is compatible with colimits which finishes the proof. ∎
From this, we deduce continuity for constructible motives.
Corollary 3.8.
Let $\Lambda$ be a commutative ring which is flat over the integers. Consider a qcqs finite-dimensional scheme $X$ that is the limit of a projective system of qcqs finite-dimensional schemes $(X_{i})$ with affine transition maps. Then, the natural map
is an equivalence.
Proof.
We first prove that the functor above is fully faithful. Let $(M_{i})$ and $(N_{i})$ be cartesian sections over the diagram of schemes $(X_{i})$ of the functor $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(-,\Lambda)$, and let $M$ and $N$ be their respective pullbacks along the maps $X\to X_{i}$, we want to prove that the canonical map
$\mathrm{colim}_{i}\hskip 2.84526pt\mathrm{Map}(M_{i},N_{i})\to\mathrm{Map}(M,N)$
is an equivalence. This can be tested after tensoring with $\mathbb{Q}$ and $\mathbb{Z}/p\mathbb{Z}$ for $p$ any prime number. Thus applying 3.1, we can assume $\Lambda$ to be a $\mathbb{Q}$-algebra or a $\mathbb{Z}/p\mathbb{Z}$-algebra. In the case of a $\mathbb{Q}$-algebra, this then follows from 3.7 and 3.2 and in the case of a $\mathbb{Z}/p\mathbb{Z}$-algebra, this follows from rigidity [BM21, Theorem 2.2].
We now prove the essential surjectivity of the functor. We use the following result.
@@
We now finish the proof. Let $M$ be a constructible motive. We want to prove that $M$ is in the essential image of the functor
$\mathrm{colim}_{i}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X_{i},\Lambda)\to\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda).$
Write the exact triangle $P\to M\oplus M[1]\to C$ of the above lemma. 3.2 implies that $P$ belongs to the essential image. As constructible motives are idempotent-complete, we can assume that $M$ is of the form $C/n$. But then, this follows from continuity with torsion coefficients which is a combination of the rigidity theorem [BH21, Theorem 3.1] and continuity for constructible étale sheaves [BM21, Theorem 2.2]. ∎
@@
Let $\Lambda$ be a commutative ring and let $X$ be a qcqs finite-dimensional scheme. Then
@@
Let $\Lambda$ be a ring which is flat over the integers and regular. Let $k$ be a field of characteristic exponent $p$. Then, the Artin motive functor
$\rho_{!}\colon\mathrm{D}(k_{\mathrm{\acute{e}t}},\Lambda[1/p])\to\mathrm{DM}^{\mathrm{\acute{e}t}}(k,\Lambda)$
@@
This follows from [Rui22, Corollary 1.2.3.4, Corollary 2.2.2.7]. ∎

### v2

@@ Abstract
Consider a scheme $X$ and a derived [^1] We define our $\infty$-categories with derived coefficients so that we may use derived rings of the form $\Lambda\otimes_{\mathbb{Z}}^{L}\mathbb{Z}/n\mathbb{Z}$. However, all main theorems are stated for ordinary commutative rings as the extension to derived rings with heavy boundedness assumptions seems artificial. commutative ring of coefficients $\Lambda$. We denote by $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$ the stable $\infty$-category of étale motives defined as the Tate-stabilisation of the category of $\mathbb{A}^{1}$-local étale hypersheaves of $\Lambda$-modules over the category of smooth $S$-schemes. This version of étale motives without transfers was studied by Ayoub in [Ayo14]; it is equivalent over Noetherian schemes to the category of h-motives studied by Cisinski and Déglise in [CD16]. It is endowed with the six functors formalism by [AGV22, Theorem 4.6.1]. Denote by $\Lambda_{X}$ the unit object of $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$.
@@
An ind-regular ring is a filtered colimit of regular rings. Our main goal is to prove that motives of geometric origin form an étale hypersheaf and to give a purely categorical characterisation of geometric étale motives:
@@
(4.1 and 2.5) Let $\Lambda$ be a commutative ind-regular ring.
@@
This question dates back to [CD16], where Cisinski and Déglise considered étale-locally geometric [^2] Beware that in loc. cit., they call constructible what we chose to call of geometric origin in this paper. We chose this terminology because they are the motives that can be built from objects coming from geometry. In addition, objects that are called locally constructible by Cisinski and Déglise are more similar to classical constructible sheaves. In any case, the purpose of this paper is to show that all those notions coincide. motives and showed that over Noetherian schemes, the latter coincides with constructible motives. This result does not hold for a general ring of coefficients (see 4.5).
To prove the above theorem, we need to generalise to non-Noetherian schemes the continuity properties proved in loc. cit.
@@
(Continuity; 3.5) Let $\Lambda$ be a commutative ring. Consider a qcqs finite-dimensional scheme $X$ that is the limit of a projective system of qcqs finite-dimensional schemes $(X_{i})$ with affine transition maps. Then, the natural map
$\mathrm{colim}_{i}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X_{i},\Lambda)\to\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda)$
@@
Using continuity, the most critical case of the main theorem is the case where $X=\mathrm{Spec}(k)$ with $k$ a field and where $\Lambda$ is regular. In that particular case, the result can be deduced from the case of rational coefficients which was proved in [CD16] combined with results of [Rui22] on Artin motives. The latter are consequences of results of [BG23] which imply that étale sheaves of the form $f_{\sharp}\Lambda_{L}$ for $L/k$ a finite separable extension, generate the category of étale-locally perfect constant sheaves as a thick subcategory of itself.
@@
(5.2 and 5.3) Let $\Lambda$ be a commutative ring (resp. an ind-regular commutative ring). The functor $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(-,\Lambda)$ satisfies Milnor excision (resp. h-descent) over qcqs finite-dimensional schemes.
@@
Let $X$ be a scheme and let $\Lambda$ be a derived [^3] As mentioned above, we make the definition here for a derived ring so that we may use it for torsion rings of the form $\Lambda\otimes^{L}_{\mathbb{Z}}\mathbb{Z}/n\mathbb{Z}$ with $\Lambda$ a commutative ring. ring. We denote by $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$ the subcategory of $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$ made of constructible motives, that is those motives such that for any open affine $U\subseteq X$, there is a finite stratification $U_{i}\subseteq U$ made of constructible locally closed subschemes such that each $M|_{U_{i}}$ is dualisable.
Remark 2.2.
If $X$ is Noetherian, then if $K$ is an object of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$, we can find a finite stratification $(X_{i})$ of $X$ such that each $K_{\mid X_{i}}$ is dualisable.
@@
For $\Lambda$ a derived ring, denote by $\mathrm{Sh}(X_{\mathrm{\acute{e}t}},\Lambda)$ the $\infty$-category of étale hypersheaves of $\Lambda$-modules. The above proposition gives a control on the rationalised part of our categories, and the following theorem will be the control on the torsion part.
Theorem 3.2 (Bachmann).
Let $n\in\mathbb{N}^{*}$ be an integer, let $\Lambda$ be a derived $\mathbb{Z}/n\mathbb{Z}$-algebra and let $X$ be a qcqs scheme such that $n$ is invertible on $X$. The natural functor
$\rho_{!}\colon\mathrm{Sh}(X_{\mathrm{\acute{e}t}},\Lambda)\to\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$
is an equivalence. We will denote by $\mathrm{D}_{\mathrm{cons}}(X_{\mathrm{\acute{e}t}},\Lambda)$ the inverse image of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda)$ under this equivalence: as $\rho_{!}$ is symmetric monoidal it has the same description as 2.1, we call those constructible étale sheaves.
Proof.
Thanks to [BH21, Theorem 3.1] the above functor is an equivalence if $\Lambda=\mathbb{Z}/n\mathbb{Z}$, and then this form of the theorem follows by taking $\Lambda$-modules on both sides. ∎
Lemma 3.3 (Martini-Wolf [MW22, Corollary 7.4.12]).
Let $\Lambda$ be a derived ring and let $X$ be a qcqs scheme. An object $K$ of $\mathrm{Sh}(X_{\mathrm{\acute{e}t}},\Lambda)$ is dualisable if and only if there exists an étale covering $f\colon U\to X$ such that $f^{*}K$ is the constant sheaf associated to a perfect complex.
Proposition 3.4.
Let $\Lambda$ be a commutative ring, let $R$ be a perfect derived $\Lambda$-algebra and let $X$ be a qcqs scheme. Assume that $X=\lim_{i\in I}X_{i}$ is a limit of qcqs schemes with affine transitions. Then the natural functor
$\mathrm{colim}_{i}\mathrm{D}_{\mathrm{cons}}((X_{i})_{\mathrm{\acute{e}t}},R)\to\mathrm{D}_{\mathrm{cons}}(X_{\mathrm{\acute{e}t}},R)$
is an equivalence, where the colimit is taken in $\mathrm{Cat}_{\infty}$.
Proof.
This is an adaptation of [BM21, Proposition 5.10]. Up to writing each $X_{i}$ as a limit of Noetherian schemes with affine transition, we can assume the $X_{i}$ to be Noetherian. We first show that the functor is fully faithful. Let $i_{0}\in I$. It is enough to prove that for any two complexes $K$ and $L$ in $\mathrm{D}_{\mathrm{cons}}((X_{i_{0}})_{\mathrm{\acute{e}t}},R)$ the map
$\mathrm{colim}_{i\geqslant i_{0}}\mathrm{Map}_{\mathrm{D}_{\mathrm{cons}}((X_{i})_{\mathrm{\acute{e}t}},R)}(K_{\mid X_{i}},L_{\mid X_{i}})\to\mathrm{Map}_{\mathrm{D}_{\mathrm{cons}}(X,\Lambda)}(K_{\mid X},L_{\mathrm{\mid X}})$
is an equivalence. By dévissage and localisation, we may assume that $K$ is dualisable and then reduce to the case where $K=R$. By adjunction, we thus have to show that the map
$\mathrm{colim}_{i\geqslant i_{0}}\mathrm{Map}_{\mathrm{Mod}_{\Lambda}}(\Lambda,\mathrm{R}\Gamma(X_{i},L))\to\mathrm{Map}_{\mathrm{Mod}_{\Lambda}}(\Lambda,\mathrm{R}\Gamma(X,L))$
is an equivalence. Now, as perfect $R$-modules are perfect $\Lambda$-modules because $R$ is perfect over $\Lambda$, the object $L$, seen as an object of $\mathrm{Sh}((X_{0})_{\mathrm{\acute{e}t}},\Lambda)$ is constructible, in particular bounded so that up to a dévissage, it is concentrated in one degree. The result then follows from the commutation of étale cohomology with colimits over the $(X_{i})$.
For the essential surjectivity, by dévissage it is enough to prove that any dualisable object lies in the image. By 3.3, if $K$ is a dualisable object of $\mathrm{Sh}(X_{\mathrm{\acute{e}t}},R)$ there exists an étale covering $f\colon U\to X$ and a perfect complex of $R$-modules $P$ such that $f^{*}(K)$ is the constant sheaf with value $P$. By Zariski descent we may even assume that $X$ is affine and up to cover the source of $f$ by an affine Zariski cover, we may also assume that the source of $f$ is affine. Then by spreading out ([EGA IV, Théorème 8.10.5, Proposition 17.7.8]) the morphism $f$ is the pullback of some étale $f^{i}$ with target $X_{i}$. Let $(f_{n})$ and $(f^{i}_{n})$ be the Čech nerves of $f$ and $f^{i}$. As the image of $(f_{n}^{i})_{\sharp}P$ in $\mathrm{D}_{\mathrm{cons}}(X,R)$ is $(f_{n})_{\sharp}P$, fully faithfulness ensures that the simplicial object $((f_{n})_{\sharp}P)_{n\in\Delta}$ is the image of a simplicial object $((f_{n}^{i})_{\sharp}P)_{n\in\Delta}$ in $\mathrm{D}_{\mathrm{cons}}(X_{i},R)$. By étale descent, we have that $K=\mathrm{colim}_{n}(f_{n})_{\sharp}P$ in $\mathrm{D}_{\mathrm{cons}}(X,R)$. Moreover if we denote by $L=\mathrm{colim}_{n}(f_{n}^{i})_{\sharp}P$ in $\mathrm{Sh}((X_{i})_{\mathrm{\acute{e}t}},R)$, we have that $(f^{i})^{*}L=P$ hence $L$ is dualisable. As the pullback functor $p_{i}^{*}\colon\mathrm{Sh}((X_{i})_{\mathrm{\acute{e}t}},R)\to\mathrm{Sh}(X_{\mathrm{\acute{e}t}},R)$ commutes with colimits, we obtain that $K$ is the image of $L$, which is a constructible object over $X_{i}$, finishing the proof.
@@
We can again assume that each $X_{i}$ is Noetherian. Fully faithfulness can be checked after tensorisation of the mapping spectra by $\mathbb{Q}$ and by $\mathbb{Z}/p\mathbb{Z}$. By 3.1, the $-\otimes\mathbb{Q}$ passes inside the mapping spectra, and then our objects become compact by [EHIK21b, Lemma 5.1 (i)] and this follows from continuity of the presentable category as in [AGV22, Lemma 2.5.11]. Modding out by $p$ passes in the mapping spectra; hence, noting that $\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z}$ is a perfect $\Lambda$-algebra, the result will then follow from the rigidity theorem in the form of 3.2 and 3.4, once we know that geometric objects are constructible with torsion coefficients. This last assertion is true over $\mathbb{Z}/p\mathbb{Z}$ by [CD16, Theorem 6.3.26]; as the functor $-\otimes_{\mathbb{Z}}\Lambda$ preserves constructible étale sheaves and as the generators of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z})$ are of the form $M\otimes_{\mathbb{Z}}\Lambda$ for some $M$ in $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(X,\mathbb{Z}/p\mathbb{Z})$, the result is true with $\Lambda\otimes_{\mathbb{Z}}\mathbb{Z}/p\mathbb{Z}$-coefficients.
To prove the essential surjectivity, first using that the Čech nerve of a covering coming from a finite open covering is finite, we can reduce to $X$ affine using Zariski descent. Then Zariski descent again (but this time on the functor that sends $g\colon Y\to X$ smooth to $g_{\sharp}g^{*}\Lambda_{X}$) implies that it suffices to reach $f_{\sharp}(\Lambda_{Y})(n)$ with $Y$ and $X$ affine and $f\colon Y\to X$ smooth. The usual spreading out properties of [EGA IV, Théorème 8.10.5, Proposition 17.7.8] then give the result. ∎
@@
For the first point, assume first that $f=p$ is proper. Let $M$ be an object of $\mathrm{DM}^{\mathrm{\acute{e}t}}_{\mathrm{gm}}(Y,\Lambda)$, and write $X=\lim X_{i}$ as a limit of schemes that are of finite type over the integers. By [EGA IV, Théorème 8.10.5.(xii)] one can find a $p_{i_{0}}\colon Y_{i_{0}}\to X_{i_{0}}$ proper such that the pullback to $X$ is $p$. In particular
@@
such that $F(j)$ is of the form $p^{j}_{*}\Lambda(n_{j})[m_{j}]$ for $p^{j}$ projective and $m_{j}$ and $n_{j}$ integers, and such that $M_{i}$ is a direct factor of $\mathrm{colim}_{J}F$. Thus, the object $M$ is a direct factor of $(\mathrm{colim}_{J}F)_{\mid X}$, which by proper base change is the colimit of the objects $q^{j}_{*}\Lambda(n_{j})[m_{j}]$, with $q^{j}$ the base change of $p^{j}$ to $X$. This finishes the proof, as the third point readily follows from the second one. ∎
As a consequence of continuity, we see that geometric motives are constructible.
@@
Let $\Lambda$ be a commutative ring. Geometric motives with coefficients $\Lambda$ are constructible.
Proof.
Let $X$ a qcqs scheme. Write $X$ as a limit of Noetherian finite-dimensional schemes $X_{i}$ with affine transition maps. Any geometric motive with coefficients $\mathbb{Z}$, over $X_{i}$ is constructible using [CD16, Theorem 6.3.26] then this is also true with coefficients $\Lambda$ by tensoring by $\Lambda$ over $\mathbb{Z}$. If $M$ is a geometric object, it is the pullback of some $M_{i}$ over some $X_{i}$. As the motive $M_{i}$ is constructible, so is $M$. ∎
We now prove the converse when $X$ is $\Lambda$-finite [^4] See the beginning of [EHIK21b, Section 5] for the precise definition. This hypothesis implies that $\mathrm{DM}^{\mathrm{\acute{e}t}}(X,\Lambda)$ is compactly generated and any finite dimensional scheme is $\Lambda$-finite if $\Lambda$ is a $\mathbb{Q}$-algebra. by linking both notions to compactness:
@@
We proved in 3.7 that motives of geometric origin are constructible. We now prove the converse. Let $M$ be constructible over $X$. We can assume that $X$ is affine by Zariski hyperdescent of geometric motives. By dévissage and localisation, we can further assume that $M$ is dualisable. In that case $\underline{\mathrm{Hom}}(M,-)$ is compatible with colimits which finishes the proof. ∎
From this, we deduce continuity for constructible motives. We begin with a general lemma:
@@
Corollary 3.10.
Let $\Lambda$ be a commutative ring. Consider a qcqs finite-dimensional scheme $X$ that is the limit of a projective system of qcqs finite-dimensional schemes $(X_{i})$ with affine transition maps. Then, the natural map
is an equivalence.
Proof.
We first prove that the functor above is fully faithful. This can be checked by tensoring the mapping spectra with $\mathbb{Q}$ and $\mathbb{Z}/p\mathbb{Z}$, and then it follows from 3.1, 3.8 and 3.4.
We now prove the essential surjectivity of the functor. Let $M$ be a constructible motive. We want to prove that $M$ is in the essential image of the functor
$\mathrm{colim}\hskip 2.84526pt\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X_{i},\Lambda)\to\mathrm{DM}^{\mathrm{\acute{e}t}}_{c}(X,\Lambda).$
Write the exact triangle $P\to M\oplus M[1]\to C$ of the above lemma. 3.5 implies that $P$ belongs to the essential image. As constructible motives are idempotent-complete, we can assume that $M$ is of the form $C/n$. But then, this follows from continuity with torsion coefficients which is a combination of 3.2 and 3.4. ∎
@@
Let $\Lambda$ be a commutative ind-regular ring and let $X$ be a qcqs finite-dimensional scheme. Then
@@
Let $\Lambda$ be a commutative ring. Let $k$ be a field of characteristic exponent $p$. Then, the Artin motive functor
$\rho_{!}\colon\mathrm{D}(k_{\mathrm{\acute{e}t}},\Lambda[1/p])\to\mathrm{DM}^{\mathrm{\acute{e}t}}(k,\Lambda[1/p])\xleftarrow{\sim}\mathrm{DM}^{\mathrm{\acute{e}t}}(k,\Lambda)$
@@
If the result is known for $\Lambda=\mathbb{Z}$, then the result for an arbitrary ring of coefficients $\Lambda$ follows from tensoring both sides with $\mathrm{Mod}_{\Lambda}$: this operation preserves full faithfulness according to [Hai22, Lemma 2.14]. The result then follows from [Rui22, Corollary 1.2.3.4, Corollary 2.2.2.7].∎